OpenStudy (anonymous):

I have been stuck on this for hours PLEASE ANYONE the length of each side of a square is 3 inches more than the length of each side of a smaller square. The sum of the areas of the squares is 149square inches.Find the lengths of the sides of the two squares.

OpenStudy (anonymous):

Let x be the length of the side of the small square. The large square has length x+3. x^2 + (x+3)^2 = 149

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

x^2+x^2+6x+9=149 2x^2+6x=140 x^2+3x=70 x^2+3x-70=0

OpenStudy (anonymous):

(x+10)(x-7)=0 so x = -10, 7

OpenStudy (anonymous):

It can't be -10 (as the length can't be negative) so the lengths must be 7 and 10

OpenStudy (anonymous):

Solving the last one you get\[(x+10)(x-7)=0\] or \[x=7\] for the smaller side

OpenStudy (anonymous):

thank you not sure where i went wrong